Exercise 4.31 x Your answer is incorrect. Try again. The Cullumber packages and distributes three grades of animal feed. Theanswers in red are wrong.
Answer #1

Fixed cost is the cost which is allocated equally among all the manufactured units.

so here given fixed cost is 10500 on total = 9400 units (1900+2800+4700).

Increase in 900 units increase fixed cost by 1800.

hence total fixed cost is 10500+1800= 12300

total no. of units= 9400+900=10300

so, fixed cost per unit is 12300/10300 = 1.19

fixed cost of 900 units = 1.19*900 =1071

variable cost of 900 units= 8*900= 7200 (premium category)

total cost for 900 units = 1071+7200= 8271

marked up cost is 200% = 8271*200%= 16542